caydenisdead:

x=? +12= 18 solve for x

KingAj:

@supie

caydenisdead:

x= 6

4evablood:

are there answer choices

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @caydenisdead x= 6 \(\color{#0cbb34}{\text{End of Quote}}\) Do you know how to get the answer?

noahalvarez:

i will hep you

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @caydenisdead x= 6 \(\color{#0cbb34}{\text{End of Quote}}\) no direct answers

umbaba5:

6+12 = 18

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Cupcake123456 \(\color{#0cbb34}{\text{Originally Posted by}}\) @caydenisdead x= 6 \(\color{#0cbb34}{\text{End of Quote}}\) no direct answers \(\color{#0cbb34}{\text{End of Quote}}\) they asked the question :/

xxDeppressionxx:

Please refrain from giving a direct answer, Thank you

xxDeppressionxx:

if you try

Cupcake123456:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @Cupcake123456 \(\color{#0cbb34}{\text{Originally Posted by}}\) @caydenisdead x= 6 \(\color{#0cbb34}{\text{End of Quote}}\) no direct answers \(\color{#0cbb34}{\text{End of Quote}}\) they asked the question :/ \(\color{#0cbb34}{\text{End of Quote}}\) x=6 isnt a question

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker \(\color{#0cbb34}{\text{Originally Posted by}}\) @caydenisdead x= 6 \(\color{#0cbb34}{\text{End of Quote}}\) Do you know how to get the answer? \(\color{#0cbb34}{\text{End of Quote}}\) X equals one so 1x - 12 = 18 then you add 12 to each side 1x = 18 + 12 so you get 1x = 30 divide both sides by 1 to get x, x = 30 hope this helps!! 😁😁